The weight of brains from Alzheimer cadavers varies according to a Normal distribution with mean 1077g and standard deviation 106g. The weight of an Alzheimer-free brain averages 1250 g. What proportion of brains with Alzheimer disease will weigh more than 1250 g?

a.

84.94%

b.

5.16%

c.

94.84%

d.

1.63%
Answer #1

Solution:

Given in the question

Mean = 1077

Standard deviation = 106

We need to calculate P(X>1250) = 1-P(X<=1250)

Z = (1250-1077)/106 = 1.63

From Z table we found p-value

P(X>1250) = 1- 0.9484 = 0.0516

So it's answer is B. I.e. 5.16%